OpenStudy (anonymous):

Please show all of your work. The area of the top of a triangular corner desk must be at least 7 sq. ft. If the base of the triangular desk top needs to be four feet, what should the height be?

OpenStudy (anonymous):

area for a triangle is (b*h)/2, so you have the value for the base, and the area, find h

OpenStudy (anonymous):

Area=(1/2)base*height 7=(1/2)(4)*height 7=2(height) height=7/2